OpenStudy (anonymous):

Hey guys, this is a seemingly easy trig question but I cant seem to get the right answer to sumbit online. The question is (Cosx)^2-1.6cosx-0.57=0. Solve on the interval [0,2pi]. So I did the quadratic and I get x=1.9 and x=-0.3. I have tryed a bunch of steps from here but none are getting the answer. Any insight would be greatly appreciated, thanks!

OpenStudy (loser66):

you have 2 solutions right? cos x = 1.9 and cos x = -0.3 Pay attention, your quadratic is cos^2 ..... not x ^2. therefore, the solutions must be respect to cos x reject the first value, because cos of an angle cannot exceed 1 So, you just have cos x = -0.3 which leads to x = 1.875

OpenStudy (loser66):

let me try, I forgot which mode I set up on my Ti 83, so, not sure that value is in radian or degree

OpenStudy (jdoe0001):

http://www.felderbooks.com/papers/ti.html

OpenStudy (loser66):

ok, it says x = 107.46 degree.

OpenStudy (loser66):

thank you mathematician jdoe0001 :)

OpenStudy (anonymous):

Thanks a lot for the help, and yup Ive tried the 107.46, it is not the right answer.

OpenStudy (loser66):

ok

OpenStudy (loser66):

I am useless here. hehehe... @jdoe0001

OpenStudy (jdoe0001):

hmm, lemme do a quickie

OpenStudy (jdoe0001):

\(\bf cos^2(x)-1.6cos(x)-0.57=0\\ \textit{let's multiply both sides by 10}\\ 10cos^2(x)-16cos(x)-57=0\\ \quad \\ x = \cfrac{ - b \pm \sqrt { b^2 -4ac}}{2a}\\ \quad \\\implies cos(x) = \cfrac{ - (-16) \pm \sqrt { (-16)^2 -4(10)(-57)}}{2(10)}\\ \quad \\ cos(x) = \cfrac{ 16 \pm \sqrt { 256 +2280}}{20}\implies cos(x) = \cfrac{ 16 \pm \sqrt { 2536}}{20}\\ \quad \\ cos(x) \implies \begin{cases} \cfrac{16+50.35871325}{20} \approx 3.32\\ \quad \\ \bf \cfrac{16-50.35871325}{20} \approx 1.72 \end{cases}\)

OpenStudy (jdoe0001):

well.. negative on the last one... \(\bf cos(x) \implies \begin{cases} \cfrac{16+50.35871325}{20} \approx 3.32\\ \quad \\ \bf \cfrac{16-50.35871325}{20} \approx -1.72 \end{cases}\)